OpenStudy (anonymous):

A charitable fund's balance is modeled by y = -4.1x2 + 7.1x + 145, where y is the balance and x is the years after 1996. If no money is added to the fund, in how many years will the fund have a balance of 0? Round to the nearest year.

OpenStudy (campbell_st):

so you have 0 = -4.1x^2 + 7.1x + 145 solve using the\[x = (-7.1 \pm \sqrt{7.1^2 - 4 \times -4.1\times145})/(2\times-4.1)\] general quadratic formula a = -4.1 b = 7.1 and c = 145 should be enough information to solve the problem